OpenStudy (anonymous):

Graph the line with slope -2/5 that passes through (3, 0).

OpenStudy (anonymous):

Use this : Y2-Y=m(X2-X) m=-2/5 X2=3 Y2=0 Got it ?:)

OpenStudy (anonymous):

thank you :)